OpenStudy (anonymous):

How do you find the discriminant for 10n(squared)+6n+1=0 ? and then state the number and type of solutions?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

The discriminant is defined as b^2 - 4ac. your equation follows the form ax^2 + bx +c, so b^2 - 4ac = 6^2 - (4*10*1) = 36 - 40 = -4, the type of the solution is imaginary, and the number of solutions is 1, 2i is the answer

OpenStudy (anonymous):

Thank you very much, but how is 2 the answer ? wouldnt -4 be the answer?

OpenStudy (anonymous):

you are right -4 is the answer, i though i read to find root

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:) np